# Women are less comfortable expressing opinions online than men and report heightened fears for safety: Surveying gender differences in experiences of online harms

### v1

@@ 3.3.3 Behavioural effects of fears about receiving online harms
While both men and women most typically indicated they were less likely to share opinions, photos and general content as a result of fearing being targeted by online harms, these behavioural effects were significantly more pronounced for women. Logistic regression analyses showed that women were 104% more likely to say they shared opinions less, 78% more likely to say they shared photos less, and 51% more likely to say they shared content less in general than men (all ps <<<.001). There was no difference between men and women in whether they were less likely to use social media overall, p=.107). Additionally, women were 30% less likely than men to say their fears do not affect their behaviour online (p=.015).
@@ 3.4 Safety tools
We used logistic regression analyses to compare the use of each of the seven safety tools (Yes=1, No=0) between men and women. Women consistently express significantly higher use of all safety tools included. Women were 29% more likely to have disabled location sharing than men (p=.034); 26% more likely to have disabled airdrop/bluetooth than men (p=.019); 94% more likely to have made an account private (p<<<.001); 55% more likely to have limited who can contact then (p<<<.001); 50% more likely to have limited who can engage with their content (p<<<.001); 60% more likely to have limited who can tag them in posts (p<<<.001); and 34% more likely to have limited how people can find them online (p=.002). See table 3 for proportions of women and men using each of the seven safety tools.
@@
Disabled airdrop and/or Bluetooth Female Yes 72.8%
@@
Made profile/account/ page private Female Yes 88.9%
@@
Limited who can contact you Female Yes 84.2%
@@
Limited who can engage with your content Female Yes 76.7%
@@
Limited who can tag/mention you in posts Female Yes 70.3%
@@
Limited how people can find your profile Female Yes 70.5%
@@ 3.5 Comfort with online behaviours
Just 23% of women are comfortable expressing political opinions online compared to almost 40% of men, 50% of women are comfortable expressing other opinions online compared to 64% of men, 22% of women are comfortable challenging content they disagree with online compared to 40% of men, and 7.5% of women are comfortable sharing personal information online compared to 11% of men. These results mean that women are 51% less comfortable than men expressing political opinions online, 44% less comfortable than men expressing other opinions online, and 57% less comfortable than men challenging content they disagree with online (all ps<<<.001). Women are also 34% less comfortable sharing personal information online (p=.008). There were no gender differences in comfort levels for sharing photos of self (23% of women and men are comfortable), of friends and family (19% of women and 18% of men are comfortable), and of activities (32% of women and 33% of men are comfortable) (all ps>>>.05). Figure 3 shows self-reported comfort with the seven online behaviours.
Figure 3: Comfort with seven online behaviours by gender differences. *** indicates a significant difference at p<<<.001; ** indicates a significant difference at p<<<.01.
@@ 3.6 Associations between fear of harm and comfort with online behaviours
To test for relationships between fears about online harms and comfort engaging in several forms of online participation, we ran a series of 84 logistic regression analyses, each including fear as predictor (one for fear about each of 12 harms), and comfort engaging in the online behaviour of interest as the outcome (one for comfort with each of the seven behaviours), all controlling for gender, age, and overall internet use. Instead of including a regression model for all fifteen harms, we excluded fear of receiving misinformation, catfishing, and eating disorder content, where there is less theoretical basis to suppose an association with comfort with the seven behaviours of interest. [^6] The remaining 12 harms we included are mostly considered ‘contact’ harms where there is risk of being directly targeted by harmful behaviours. Hate speech and misogyny may be considered ‘content’ harms in general, but still pose risk as contact harms if respondents’ social identities are targeted. In each model, we used a binary predictor for fear of each harm and a binary outcome for comfort with each behaviour, as described in the relevant sections above.
@@
On average, being fearful of receiving online harms is associated with being 25% less comfortable with sharing opinions, photos or other information online. Furthermore, fear of different harms impacts on comfort participating with different online behaviours. For example, fear of being targeted by trolling is associated with reduced comfort with sharing political opinions, other opinions, and challenging content. Fear of being targeted by misogyny, cyberstalking and doxing is associated with reduced comfort with sharing photos. Full results for each regression model can be found in LABEL:tab:Regressions in Supplementary Information.
Figure 4: Fear of directly receiving harms reduces comfort with online behaviours. Here, significant associations between fears and comfort with behaviours are shown in different coloured (and shaped) points for each online behaviour. The black line highlights the average percentage change in comfort with online participation associated with fear of harm, and the red line highlights the average percentage change in comfort with online participation associated with gender. Note sharing personal information is not included as none of these models were significant.
@@ 4 Discussion
Using a large, nationally representative survey of UK adults, we examined gender differences in exposure to harmful content online, along with the impact of online experiences. We examined the extent to which people had been exposed to fifteen different types of online harms, fears surrounding such exposure, the psychological impact of online experiences, the use of safety tools online, and comfort with several forms of online participation.
@@
When asked about whether they had seen different forms of harmful online content over the last year, men and women reported similar levels of exposure overall, consistent with work asking broadly about how much people have seen content which they consider to be harmful online (Enock et al.,, 2023). However, by asking about fifteen specific harms and the capacity in which people have seen them, we were able to detect more nuance in experiences. Here, women report being directly targeted by certain forms of of ‘contact-based’ harms like cyberflashing, cyberstalking and image based abuse to a greater extent than men, consistent with related prior work (Storry and Poppleton,, 2022; United Nations,, 2018; Amnesty International,, 2017). However, men report being directly targeted by threats of physical violence, hate speech and trolling to a greater extent than women. These results are more novel and further work should examine why men may be more at risk of certain harms. Our findings show that our approach in including men and women in our sample and asking about exposure to a range of online harms is beneficial in gaining a more nuanced picture of gender differences in exposure to online harms.
@@
As well as being targeted by harms such as cyberflashing and image based abuse to a greater extent than men, we found that women report having experienced greater negative psychological impact than men as a result of certain experiences online, including feeling sad and low, angry and frustrated, and with physical symptoms such as insomnia or headaches. Our results extend prior work showing that women may be more negatively affected by online harms than men (Storry and Poppleton,, 2022) by demonstrating increased negative affect in a representative sample of women, not only those who have been direct targets of abuse or harassment. Our results are important in contributing to our understanding of the ways in which women are disproportionately affected by online experiences and in how support should best be directed. If women are more likely to feel sad, frustrated or physically unwell after online experiences, it is crucial that we do more to protect their online wellbeing.
@@
Consistent with our finding that women report more negative psychological impact as a result of online experiences than men, we found that women expressed greater fear of being targeted by all fifteen harms that we asked about compared to men. These results show that women carry an additional psychological burden relating to safety concerns online. While no work to our knowledge had examined gender differences in fears about experiencing different forms of online harm, our findings accord with work in the offline crime literature showing that women typically fear for their personal safety more than men do (Snedker,, 2012). We show these additional fears extend to the online world. Both men and women indicated that fear of exposure to online harms most typically originate from the media and from female friends’ experiences, suggesting that along with the role of the media in perpetuating safety concerns (consistent with Cashmore,, 2014), the known negative experiences of other women are also highly salient.
@@
It is interesting to consider the interplay between fear and exposure. While women consistently express greater fear about being targeted by harms, they are not always at greater risk of such exposure. It is important to question whether women’s fears are greater because the types of harm they are at greater risk of being targeted by are particularly threatening (e.g., image based abuse), or whether increased fears lead women to do more to protect themselves online in the form of safety work, thereby protecting them somewhat from increased exposure in the future. As we analyse only self-report data, we also cannot rule out the fact that women may be more comfortable expressing fear than men are.
In line with our conjecture that women may do more safety work to protect themselves from harmful online experiences than men, we found that women report using all seven safety tools that we asked about to a greater extent than men. This result extends recent findings from Bright et al., (2024), showing that across a range of safety tools and comparing men and women in a nationally representative sample (not just social media users who have already heard of the tool), women are engaging in more safety work to protect themselves online. This additional safety work may be driven by fear, by prior exposure to harm, or both.
@@
As well as engaging with safety tools to a great extent than men, women also report being less comfortable than men with participating in several online behaviours including expressing political opinions online, expressing opinions more generally, and challenging content that they disagree with. While some prior work shows that women conduct themselves differently online after experiencing abuse (Amnesty International,, 2017), our findings are novel in directly comparing men and women’s general comfort with online participation. Our results suggest that women may be self-censoring online, particularly in terms of expressing opinions and challenging other content. These results are concerning, because with much public discourse happening online, gender inequality in public spaces is likely to be perpetuated if women feel too fearful to participate.
@@
It is plausible that greater fear surrounding exposure to online harms plays a causal role in women’s self-censorship online. Indeed, we found that women were more likely than men to say that they are less likely to share opinions, photos and general content as a direct result of fearing being targeted by online harms. These patterns accord with offline behaviours whereby fear of crime leads to moderation of one’s actions, such as avoiding walking in quiet places after dark, or taking extra precautions when travelling alone (Office for National Statistics,, 2022). We directly tested associations between fears about being targeted by twelve online harms and comfort with the seven online behaviours included in the survey. Overall, being more fearful was associated with being less comfortable with several forms of online participation, and different fears influenced comfort with different online behaviours when gender, age and internet use was controlled for. For example, fear of directly receiving online misogyny significantly predicted less comfort expressing political opinions online, while fear of cyberstalking significantly predicted less comfort sharing photos of the self and of friends and family online. It may be that greater exposure to particular forms of harm increase fear of being targeted by harms more generally in women, leading to additional safety work in the form of using safety tools and reducing online participation, perhaps mitigating risk of further exposure. Alternatively, it may be that greater fears amongst women arise by factors other than direct exposure, leading to increased safety work, and again mitigating further risk. Our work shows that the interplay between gender, exposure to harm, fears and online behaviours is complex and future work would benefit from examining these links in greater detail.
@@
While our work offers important novel insights into the differential impacts of online harms experienced by men and women, it is important to acknowledge limitations in the study and outstanding research questions that were not possible to answer with this data. Firstly, in attempting to understand gender differences in experiences of online harms using a nationally representative sample, we compare responses given by people identifying as men or women, excluding individuals who identify as non-binary or another gender. The reason for taking this approach is because in a nationally representative sample of 2000 people, numbers of individuals identifying as a different gender are too small (in this sample, 17 individuals identified as either non-binary, preferred to self-describe or preferred not to say) to make meaningful statistically generalisable comparisons for this group. Future work could look more specifically at the experiences of individuals identifying in other ways by oversampling this group and actively engaging this sample. Additionally, our sample comprised only of adults, and future work would benefit from tackling similar research questions with children and teenagers, who are also at risk of gendered online harms (Plan International,, 2020; Livingstone,, 2014).
@@
In our analyses, we generalise across men and women as a whole, but there are likely to be intersections with gender and other social features such as age and race. Future work should consider these intersections in greater detail, for example in examining differing experiences of older vs. younger women and men, or exploring gender differences based on ethnicity, for example extending findings from OfCom, (2022) which suggested Black women were least likely to think that being online has a positive effect on mental health than those from other ethnic backgrounds.
@@
Overall, we provide up-to-date evidence within a nationally representative sample of UK adults about how and when women are disproportionately affected by online harms. Our findings demonstrate the myriad of harms that men and women are exposed to online, providing valuable information about who is most at risk and when, important for knowing where psychological support should most be directed and in what ways interventions against online harms may be most effective. Our work highlights the need for greater efforts from platforms to protect women users, welcoming efforts from the Online Safety Act in requiring Ofcom to develop guidance for tech companies to reduce harm to women and girls (End Violence Against Women,, 2023). To ensure an egalitarian society, we must make sure that all members of society feel safe and able to participate in online spaces.

### v2

@@ 4.3.3 Behavioural effects of fears
While both men and women most typically indicated they were less likely to share opinions, photos and general content as a result of fearing being targeted by online harms, these behavioural effects were significantly more pronounced for women. Logistic regression analyses showed that women were 104% more likely to say they shared opinions less, 78% more likely to say they shared photos less, and 51% more likely to say they shared content less in general than men (all ps <<.001). There was no difference between men and women in whether they were less likely to use social media overall, p=.107). Additionally, women were 30% less likely than men to say their fears do not affect their behaviour online (p=.015).
@@ 4.4 Safety tools
We used logistic regression analyses to compare the use of each of the seven safety tools (Yes=1, No=0) between men and women. Women consistently expressed significantly higher use of all safety tools included. Women were 29% more likely to have disabled location sharing than men (p=.034); 26% more likely to have disabled airdrop/bluetooth than men (p=.019); 94% more likely to have made an account private (p<<.001); 55% more likely to have limited who can contact then (p<<.001); 50% more likely to have limited who can engage with their content (p<<.001); 60% more likely to have limited who can tag them in posts (p<<.001); and 34% more likely to have limited how people can find them online (p=.002). See table 3 for proportions of women and men using each of the seven safety tools.
@@
Disabled airdrop
and/or Bluetooth
Female Yes 72.8%
@@
Made profile/account/
page private
Female Yes 88.9%
@@
Limited who can
contact you
Female Yes 84.2%
@@
Limited who can
engage with your content
Female Yes 76.7%
@@
Limited who can
tag/mention you in posts
Female Yes 70.3%
@@
Limited how people can
find your profile
Female Yes 70.5%
@@ 4.5 Comfort with online behaviours
Just 23% of women are comfortable expressing political opinions online compared to almost 40% of men, 50% of women are comfortable expressing other opinions online compared to 64% of men, 22% of women are comfortable challenging content they disagree with online compared to 40% of men, and 7.5% of women are comfortable sharing personal information online compared to 11% of men. Logistic regressions showed women are 51% less comfortable than men expressing political opinions online, 44% less comfortable than men expressing other opinions online, and 57% less comfortable than men challenging content they disagree with online (all ps<<.001). Women are also 34% less comfortable sharing personal information online (p=.008). There were no gender differences in comfort levels for sharing photos of self, of friends and family, and of activities (all ps>>.05). Figure 3 shows self-reported comfort levels with the seven online behaviours.
Figure 3: Comfort with seven online behaviours by gender differences. *** indicates a significant difference at p<<.001; ** indicates a significant difference at p<<.01.
@@ 4.6 Associations between fear of harm and comfort with online behaviours
To test for relationships between fears about online harms and comfort engaging in several forms of online participation, we ran a series of 84 logistic regression analyses, each including fear as predictor (one for fear relating to each of 12 harms), and comfort engaging in the online behaviour of interest as the outcome (one for comfort with each of the seven behaviours), all controlling for gender, age, and overall internet use. Instead of including a regression model for all fifteen harms, we excluded fear of receiving misinformation, catfishing, and eating disorder content, where there is less theoretical basis to suppose an association with comfort with the seven behaviours of interest. The remaining 12 harms we included are mostly considered ‘contact’ harms where there is risk of being directly targeted by harmful behaviours. Hate speech and misogyny may be considered ‘content’ harms in general, but still pose risk as contact harms if respondents’ social identities are targeted. In each model, we used a binary predictor for fear of each harm and a binary outcome for comfort with each behaviour, as described in the relevant sections above.
@@
On average, being fearful of receiving online harms is associated with being 26% less comfortable with sharing opinions, photos or other information online. Furthermore, fear of different harms impacts on comfort participating with different online behaviours. For example, fear of being targeted by trolling is associated with reduced comfort with sharing political opinions, other opinions, and challenging content. Fear of being targeted by misogyny, cyberstalking and doxing is associated with reduced comfort with sharing photos. Full results for each regression model can be found in Table LABEL:tab:Regressions in Supplementary Information.
Figure 4: Fear of directly receiving harms is associated with reduced comfort with online behaviours. Here, significant associations between fears and comfort with behaviours are shown in different coloured (and shaped) points for each online behaviour. The black line highlights the average percentage change in comfort with online participation associated with fear of harm, and the red line highlights the average percentage change in comfort with online participation associated with gender. Sharing personal information is not included as none of these models were significant.
@@ 5 Discussion
Using a nationally representative survey of UK adults, we examined the role of gender in shaping exposure to harmful online content, fears surrounding such exposure, the psychological impacts of online experiences, engagement with safety tools, and comfort with online participation.
@@
When asked about whether they had seen different forms of harmful online content over the last year, men and women reported similar levels of exposure overall, consistent with prior work asking broadly about how much people have seen general content which they consider to be harmful online (Enock et al.,, 2023). However, by asking about fifteen specific harms and the capacity in which people have seen them, we were able to detect more nuance in experiences. Here, women report being directly targeted by certain forms of ‘contact-based’ harms like cyberflashing, cyberstalking and image based abuse to a greater extent than men, consistent with related prior work (Storry and Poppleton,, 2022; UN Human Rights Council,, 2018; Amnesty International,, 2017). These findings align with feminist accounts of the ‘continuum of violence’ (Kelly,, 1988), supporting the notion that digital abuse and harassment are part of the broader spectrum of gendered violence that extends across offline and online contexts. Our findings highlight that digital spaces are not neutral, and rather reflect and reproduce existing structural inequalities.
However, we also found that men report being directly targeted by threats of physical violence, hate speech and trolling to a greater extent than women. These results are more novel and suggest that the types of harm men and women experience online reflect offline gendered norms, where women more commonly experience sexualised and silencing harms that reproduce structural subordination, while men may encounter more overt, confrontational harms. Our findings show that our approach in including men and women in our sample and asking about exposure to a range of harms is beneficial in gaining a more nuanced picture of gender differences in exposure.
@@
We also found that women report having experienced greater negative psychological impact than men as a result of certain experiences online, including feeling sad and low, angry and frustrated, and with physical symptoms such as insomnia or headaches. Our results extend prior work showing that women may be more negatively affected by online harms than men (Storry and Poppleton,, 2022) by demonstrating increased negative affect in a representative sample of women, not only those who have been direct targets of abuse or harassment. These findings align with feminist accounts of ‘symbolic violence’ Bourdieu, (1991), where routine exposure to gender-based inequality reinforces women’s subordinate position and damages psychological wellbeing.
@@
Consistent with our finding that women report more negative psychological impact as a result of online experiences than men, we found that women expressed greater fear of being targeted by all fifteen harms that we asked about compared to men. These results show that women carry an additional psychological burden relating to safety concerns online, according with recent survey work showing that women are significantly more fearful than men of becoming the subjects of harmful deepfakes, such as deepfake pornography (Sippy et al.,, 2024). Our findings also accord with work in the offline crime literature showing that women typically fear for their personal safety more than men do (Snedker,, 2012). Feminist scholarship has long suggested that women’s disproportionate fear of violence reflects their structural positioning in a gendered world (Stanko,, 1990; Kelly,, 1988). Through this lens, the heightened fears reported by women reflect not only individual risk perception, but broader patterns of gendered safety work, where women carry the burden of responsibility for protecting themselves (Vera-Gray and Kelly,, 2020). Both men and women indicated that fear of exposure to online harms most typically originate from the media and from female friends’ experiences, suggesting that along with the role of the media in perpetuating safety concerns (consistent with Cashmore,, 2014), the known negative experiences of other women are also particularly salient.
@@
It is interesting to consider the interplay between fear and exposure. While women consistently express greater fear about being targeted by harms, they are not always at greater risk of such exposure. Feminist criminology highlights that women’s fears surrounding personal harm often exceed their actual likelihood of experiencing such harm, with fear a mechanism in the structural oppression of women, functioning to regulate women’s behaviour and restrict participation (Stanko,, 1990; Kelly,, 1988). It is important to question whether increased fears lead women to do more to protect themselves online in the form of safety work. At the same time, we cannot rule out that women are simply more comfortable expressing fear than men.
In line with our conjecture that women may do more safety work to protect themselves from harmful online experiences than men, we found that women report using all seven safety tools that we asked about to a greater extent than men. These results align with theoretical claims that women engage in disproportionate self-protection work (Vera-Gray and Kelly,, 2020). Our findings suggest that the unequal patterns of safety work long observed offline are reproduced in online spaces, reinforcing the additional burden of risk management for women.
@@
As well as engaging with safety tools to a great extent than men, women also report being less comfortable than men with participating in several online behaviours including expressing political opinions online, expressing opinions more generally, and challenging content that they disagree with. While some prior work shows that women conduct themselves differently online after experiencing abuse (Amnesty International,, 2017), our findings are novel in directly comparing men and women’s general comfort with online participation. Our results suggest that women may be self-censoring online, particularly in terms of expressing opinions and challenging other content. This can be understood as a form of gendered silencing, where fears around harm exposure interact with disproportionate safety work to constrain women’s digital citizenship and public voice (Vera-Gray,, 2018; Olson Carter and LaPoe,, 2018; Chen et al.,, 2020). These results are concerning, because with much public discourse happening online, gender inequality in public spaces is likely to be perpetuated if women feel too fearful to participate.
@@
Our findings suggest that greater fear surrounding exposure to online harms plays a key role in women’s self-censorship online. Women were more likely than men to say that they are less likely to share opinions, photos and general content as a direct result of fearing being targeted by online harms. These patterns accord with offline behaviours whereby fear of crime leads to moderation of one’s actions, such as avoiding walking in quiet places after dark, or taking extra precautions when travelling alone (Office for National Statistics,, 2022). Consistent with this, our regression analysis showed that greater fear of harm exposure was associated with lower comfort with several different online behaviours, even when controlling for gender, age, and internet use. For example, fear of directly receiving online misogyny significantly predicted less comfort expressing political opinions online, while fear of cyberstalking significantly predicted less comfort sharing photos of the self and of friends and family online. The interplay between harm exposure, fear, safety work, and participation is likely to be complex. It may be that greater exposure to particular harms increases fear in women, leading to additional safety work, or that greater fears arise by factors such as cultural transmission through media and social networks. Future research would benefit from working to understand the causal nature of these relationships in more detail, which could inform how to best target intervention.
While our work offers important novel insights into the differential impacts of online harms experienced by men and women, it is important to acknowledge limitations in the study and outstanding research questions that were not possible to answer with this data. Our Prolific sample was designed to be representative of the UK population across age, gender, and ethnicity, but because Prolific recruitment is not probability-based and relies on participants filling survey slots on a first-come, first-served basis, there remains the possibility of selection bias.
@@
Additionally, in attempting to understand gendered experiences of online harms, we statistically compare responses given by people identifying as women with those identifying as men. In our sample, a small number of respondents identified as non-binary or another gender (N=17 total), but the sample size was too limited to allow for robust statistical comparisons. Future work could look more specifically at the experiences of individuals identifying in other ways by oversampling this group and actively engaging this sample to capture distinct experiences of online harms. Additionally, our sample included only adults, and future work would benefit from tackling similar research questions with children and teenagers (Plan International,, 2020; Livingstone,, 2014).
@@
In our analyses, we generalise across men and women as a whole, but there are likely to be intersections with gender and other characteristic such as age and ethnicity. Future work should consider these intersections in greater detail, for example in examining differing experiences of older vs. younger women and men, or exploring gender differences based on ethnicity, for example extending findings from Ofcom, (2022) which suggested Black women were least likely to think that being online has a positive effect on mental health than those from other ethnic backgrounds.
@@
Overall, we provide novel evidence about how and when women are disproportionately affected by online harms. Our findings demonstrate the myriad of harms that men and women are exposed to online, providing valuable information about who is most at risk and when, important for knowing where psychological support should most be directed and in what ways interventions against online harms may be most effective. Our work highlights the need for greater efforts from platforms to protect women users, welcoming efforts from the Online Safety Act in requiring Ofcom to develop guidance for tech companies to reduce harm to women and girls (End Violence Against Women,, 2023; Ofcom,, 2025), but we also emphasise that such measures must go beyond individual focus and address the systemic reproduction of inequality in digital spaces. To ensure an egalitarian society, we must make sure that all members of society feel safe and able to participate in online spaces.